What is the probability of winning the lottery in with you have to pick 6 numbers out of 1-16?
It is 1/8008
How many different arrangements can be made by four letters of the word ENGLISH?
Answer: 7P4 = 840 arrangements
Number of letters in the word ENGLISH = 7
Number of letters to take= 4
Number of arrangements:
= 7P4
= 7!/(7-4)!
= 5040/6
= 840
What is the probability of getting an odd number greater than 5?
49.999 (repeater)%
* * * * *
It is not possible to answer the question without knowing what the experiment or the event space is.
Random Sampling.
What is the probability of tossing four coins and getting tails on all four?
Assuming it is a fair coin, the probability is 1/24 = 1/16.
Compute the probability of any specific sequence of hits and misses, such as HMHHHM, by multiplying the probabilities of each throw: 5/7 for each hit ("H") and 1 - 5/7 = 2/7 for each miss ("M"). Because there are four hits (and therefore two misses), the multiplication includes four factors of 5/7 and two factors of 2/7; briefly, the product equals 5^4 * 2^2 / 7^6. There are 15 distinct sequences of four hits and two misses. Therefore the probability of getting some sequence of exactly four hits equals 15 times this previous value: 15 * 5^4 * 2^2 / 7^6 = 37500 / 117649 = about 31.87%.
What does inference mean in statistics?
It can be defined as followed. The conclusion reached on the basis of evidence.
What is the probability of drawing a 3 out of 52 cards?
The probability of drawing a 3 out of 52 cards is 4/52 or 1/13 or 0.0769 or 7.69%.
Define and describe probable cause?
Probable cause refers to the standard by which a police officer has the right to make an arrest, conduct a personal or property search, or to obtain a warrant for arrest. It is also used to refer to the standard to which a grand jury believes that a crime has been committed.
Added: Probable cause is the standard used in justifying certain police actions. For example, police need to have probable cause to believe evidence of a crime exists when making an arrest or in requesting a search warrant.
Probable Cause is more than mere suspicion but less than the amount of evidence required for conviction.
For example: A police officer may use "probable cause" to arrest someone for attempted theft when he finds someone trespassing on private property late at night wearing a stocking mask. Using this probable cause as justification for stopping the subject and, searching them, they find them in possession of burglary tools, thereby justifying the probable cause forcible stop.
What is the probability of 1 out of 2 failing?
Any time there are two options, heads or tails when you flip a coin for example, the probability is 1/2, that the result will be either one option, or the other. The expected result when the coin lands is a 1/2 probability that it will be heads, and a 1/2 probability that it will be tails. What "1 out of 2 failing" means is that for every two students that take an exam, for example, one of them will fail. Of course, it also means, that 1/2 will pass.
When is the most dangerous time to drive a car?
When you are impaired by something for example; being under the influence
How do you use the z-score to determine a normal curve?
If the Z-Score corresponds to the standard deviation, then the distribution is "normal", or Gaussian.
What is the probability of rolling a product less than 16?
The answer depens on how many dice are rolled. And since you have chosen not to divulge that critical bit of information, it is not possible to give a more useful answer.
How do you find the area of a normal distribution?
The area under a normal distribution is one since, by definition, the sum of any series of probabilities is one and, therefore, the integral (or area under the curve) of any probability distribution from negative infinity to infinity is one. However, if you take an interval of a normal distribution, its area can be anywhere between 0 and 1.
What is the probability to get a baby on saturday in a leap year?
The answer depends on where in the world you are - much more than whether or not it is a leap year. In some countries, a majority of births take place in hospitals or maternity units. Medical staff in such establishments do not like working on Saturdays and so the probability of getting a baby on a Saturday is less than 1/7.
I haven't done statistics in a while but I think it should go like this
(1/8) * (1/7) = (1/56)
Is it possible to have an experimental probability of 0?
yes. for example role a 6 sided die, will you get a 7? never. the experimental and theoretical probability is 0
What is the probability of a family having a boy after four girls?
The probability of a boy is still 0.5 no matter how many prior children there are.
1/4
Describe the central limit theorem and give an example of how it can be used in statistics?
The Central Limit Theorem (CLT) is a theorem that describes the fact that if a number of samples are taken from a population, the distribution of the means of the samples will be normal. This is true for all different distributions, whether or not the population is normal or something else. The main exception to this is that the theorem does not work particularly well if the samples are small (<30) and the original population is not distributed normally.
The sample means will be distributed with a mean equal to the population mean, and with variance equal to the variance of the population divided by the size of each sample.
This can sound confusing when first working it out, but once it makes sense it is very useful in statistics. It is the basis for confidence intervals and hypothesis testing, as well as other statistical tests.